# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v1

@@ 5.1 Extend Sample Size Through PEFT
Drawing from the insights in lemma 3, we recognize the substantial impact of increasing task-specific sample size on reducing generalization gaps. Building on this perspective, we propose an advanced PEFT strategy operative at three subnetworks that generate features in distinct spaces. This strategy synergistically leverages features derived from dual adaptation mechanisms to forge hybrid features, effectively extending the sample size for each sub-task. By adeptly fine-tuning sub-network parameters $p_{i}$ through an amalgamation of these multifaceted features, our method expands each sample’s representational breadth threefold, as elaborated through the sophisticated adaptive interactions in Fig. 2.
@@
PEFT-CL methodologies typically employ a prompt pool or introduce auxiliary parameters while preserving pre-trained weights, modifying $E_{1}^{i},E_{2}^{i},\ldots,E_{N}^{i}$ within each transformer block to influence the class token $E_{CLS}$. This generates a novel feature space that adapts to sub-tasks and mitigates catastrophic forgetting. In these methods, the predetermined task prompt pool is traditionally used to derive task-specific embeddings, selecting prompts through cosine similarity [82, 81, 36]. While effective, this paradigm incurs substantial computational overhead when intervening in the self-attention mechanism and constrains the network’s capacity for generating diverse, instance-specific adaptive interventions dynamically. To address these limitations, our proposed NTK-CL framework implements a more efficient paradigm utilizing additional trainable parameters to autonomously generates instance-specific interventions. These interventions then interact with our proposed feature space post-multi-head self-attention (MSA) module to yield task-specific embeddings. This approach not only maximizes the utilization of pre-trained knowledge but also effectively reduces the computational burden brought by intervening MSA calculations.
@@
Figure 3: The illustration depicts the fusion of multi-level features to generate three distinct embeddings per sample, increasing the optimizable sample size.
@@ 5.2 Task-Level Feature Constraints
Therefore, we propose a compromise approach. Within the context of NTK theory, the optimization of infinitely wide neural networks mirrors a Gaussian process [10, 40], yielding a locally constant NTK matrix [32, 91, 42, 12]. Given this, it is reasonable to assume that $\Phi^{*}(X_{\tau},X_{k})=\Phi_{0}(X_{\tau},X_{k})=\Phi_{1}(X_{\tau},X_{k})=\cdots=\Phi_{\infty}(X_{\tau},X_{k})$. Moreover, networks pre-trained on extensive datasets emulate the properties of infinitely wide networks [41, 83, 75], aligning with our pre-trained model. Therefore, we relax the original constraint, assuming that the pre-trained model is at this local optimum.
@@
Knowledge Retention: Effective replay of past knowledge is essential in PEFT-CL utilizing pre-trained models. For instance, [77] employs a distributional strategy that approximates and replays past data distributions for ongoing optimization, while [100] preserves distinct adapter parameters for each historical task. These strategies, while boosting knowledge retention, simultaneously escalate optimization and storage requirements, posing challenges to scalability and operational efficiency in scenarios with large datasets or extensive task sequences. To mitigate these issues, we introduce an adaptive Exponential Moving Average (EMA) mechanism tailored for efficient past knowledge preservation, illustrated in Fig. 4.
@@
Traditional EMA applications often maintain a static base model, incrementally integrating optimized weights to preserve historical data. However, this approach proves suboptimal in PEFT-CL settings due to the substantial disparities in weights across tasks. Directly preserving a large proportion of past weights can detrimentally affect the performance on current tasks, while retaining an entire model’s weights is excessively redundant. Therefore, we propose two improvements. First, we categorize the adaptation parameters responsible for generating embedding into two segments: $p^{pre}$ for historical knowledge and $p^{curr}$ for current insights. Secondly, we apply the EMA mechanism exclusively to the adaptation modules’ parameters, leaving other optimizable parameters untouched to ensure the optimization remains streamlined. Throughout the optimization of task $\tau+1$, only $p^{curr}_{\tau+1}$ is modified, while $p^{pre}_{\tau+1}$ is adaptively adjusted post-task-$\tau$ completion, employing an adaptive EMA scheme:
@@
Under this mechanism, each past task equitably contributes to constructing embeddings related to historical knowledge without compromising the current tasks’ insights, while avoiding the excessive memory overhead of storing parameters for each task, as seen in [100]. Consequently, $E_{CLS}^{S1}$, $E_{CLS}^{S2}$, and $E_{CLS}^{HAE}$ all consist of two components: $concat[f(x,p_{pre}),f(x,p_{curr})]$.
Task-Feature Dissimilarity:  [^3] Regarding why task-feature orthogonality does not impair the propagation and retention of knowledge among similar classes across different tasks, we provide further explanations in Appendix E. Based on Sect. 5.2 and insights from [16, 3], it is evident that achieving orthogonal insulation between $f^{*}(X_{\tau})$ and $f^{*}(X_{k})$ can compromise to reduce generalization gap. Therefore, relying on the prototype classifier, we propose an optimization loss. In line with [100, 77], we update the prototype classifier $\zeta$ upon completion of each task’s optimization and strictly prohibit accessing previous samples in subsequent optimizations to comply with PEFT-CL constraints. During the optimization of task $\tau$, we randomly sample $\zeta_{\tau}$ from $\zeta$ to represent $f^{*}(X_{\tau})$  [^4] Sampling from the parameter space of the prototype classifier $\zeta$, unlike approaches such as Hide-Prompt [77] and APG [73], avoids compressing past embedding distributions and adding extra training overhead. This method also eliminates the need for a replay buffer, effectively bypassing the typical constraints associated with PEFT-CL.. To initially distinguish $f^{*}(X_{\tau})$ from $f^{*}(X_{k})$, we use the InfoNCE [57] as a metric, employing $\zeta_{\tau}$ as the negative sample, while using samples $x_{\tau}$ (represented by $E_{CLS}^{HAE}$, as this is the feature used for final classification) from task $\tau$ as positive samples.
@@
To further ensure orthogonality between $f^{*}(X_{\tau})$ and $f^{*}(X_{k})$, we apply the truncated SVD method [21] to constrain the optimization of $f^{*}(X_{k})$. Specifically, we decompose $\zeta$ to obtain the orthogonal basis $\mathbf{U}$ that defines the classification (preceding feature) space. We then map $x_{\tau}$ into this space and remove the unmappable part from the original $x_{\tau}$. When the retained mappable portion is sufficiently small, the orthogonality between $x_{\tau}$ and $\zeta$ is ensured, thus maintaining orthogonality.
@@
where $\tilde{proj}(a,b)$ is the unmappable portion of $a$ in the space formed by the orthogonal basis functions decomposed from $b$.
@@ 5.3 Regularization Adjustment
In accordance with the theoretical constraints delineated in Appendix A, which advocate for the incorporation of ridge regression to ensure a well-conditioned solution, we deploy an L2 regularization [28]. As specified in Eq. 32, the regularization term is structured as $\left\|p_{\tau}-p_{\tau-1}^{*}\right\|_{2}^{2}$, targeting the parameter shifts from task $\tau-1$ to task $\tau$. Consequently, we meticulously design our regularization term to mirror this structure and temporarily retain the trainable parameters $p^{pre}$ from the preceding task. This targeted regularization is then precisely applied to the parameters of the various modules within our NTK-CL, formulated as follows:
@@
$\mathcal{L}=\mathcal{L}_{cls}+\eta\mathcal{L}_{dis}+\upsilon\mathcal{L}_{orth}+\lambda\mathcal{L}_{reg},$ (29)
@@
Testing Evaluation: Upon concluding the training regimen, the evaluation phase commences with simultaneous testing across all tasks. This phase distinctly prioritizes the synthesized hybrid adaptation features $E_{CLS}^{HAE}$ for final analysis. Through the final prototype classifier $\zeta$, these features are transformed into logits, which are subsequently aligned with the corresponding labels to deduce the test accuracy.
@@ 6 Experiments
Datasets: In our empirical analysis, we utilize a comprehensive collection of datasets, meticulously curated to rigorously evaluate the model’s adaptability within the PEFT-CL scenarios. These datasets span a broad spectrum of domains, including general object recognition, domain adaptation, animal recognition, and geospatial analysis, providing a robust framework for assessment. Below, we elucidate the significance, structure, and relevance of each dataset to our investigation:
CIFAR-100 [38]: This dataset, crucial for assessing continual learning algorithms, comprises 60,000 32x32 color images across 100 classes, with each class contributing 600 images. For PEFT-CL, it is segmented into 10 distinct tasks, each encompassing 10 classes. Images are resized to 224x224 to be compatible with the pre-trained ViT model.
ImageNet-R [81]: It enriches the original ImageNet dataset by incorporating artworks, cartoons, and interpretations for 200 classes. This dataset underscores the challenge of generalizing across highly varied visual domains. For our evaluation, the ImageNet-R is organized into 10 tasks, each containing 20 classes, with a distribution of 24,000 training images and 6,000 test images, offering a rigorous testbed for assessing model generalization in complex PEFT-CL scenarios.
ImageNet-A [25]: This challenging dataset is an extension of ImageNet, curated to evaluate model robustness against adversarial and out-of-distribution samples. It contains 7,500 images from 200 classes, each designed to be misclassified by standard models. For PEFT-CL, ImageNet-A is divided into 10 tasks, providing a stringent test of a model’s ability to handle challenging and unusual cases, thereby assessing resilience and generalization in continual learning scenarios.
DomainNet [59]: This large-scale domain adaptation dataset includes six domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch, comprising 423,506 images across 345 categories. For PEFT-CL, it is divided into 15 tasks, each containing 23 classes, to rigorously test cross-domain generalization capabilities. Our study expands beyond previous research like DAP [36], which focused solely on the Real domain, and CODA-Prompt [71], which limited its investigation to a five-task sequence within the Real domain. By encompassing all six domains in a structured 15-task framework, we provide a comprehensive assessment of domain adaptation within the CL context, establishing a new benchmark in this research area.
Moving beyond conventional datasets utilized in CL research, our study integrates a wide array of datasets from diverse domains, encompassing object recognition, domain adaptation, animal recognition, earth observation, agricultural studies, task adaptation, and healthcare diagnosis. We employ datasets such as Oxford Pets [58], EuroSAT [24], PlantVillage [30], VTAB [98], and Kvasir [61], as summarized in Table I. This comprehensive dataset selection underscores our model’s robustness, extensive generalization, and adaptability across a spectrum of visual recognition tasks, thereby reinforcing its capability to address varied domain-specific challenges.
Evaluation Metrics: Following the established benchmark protocol as outlined by [66], we evaluate the model’s effectiveness using $A_{\tau}$, which signifies the accuracy post the $\tau$-th training stage. Notably, we employ $A_{T}$—the performance metric at the termination of the final stage—and $\bar{A}=\frac{1}{T}\sum_{\tau=1}^{T}A_{\tau}$, which calculates the average accuracy over all incremental stages. These metrics are selected as the principal measures of model performance, providing a holistic view of its efficacy and stability throughout the training process.
@@
TABLE VI: Ablation study on the ViT-B/16-IN21K model, evaluating its performance across the CIFAR100 and ImageNet-R datasets. The study features a detailed breakdown of model components, denoted in each column by the inclusion (✓) of specific modules and strategies: Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid), Knowledge Retention (KR), Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). Incremental accuracies ($\bar{A}$) are reported to highlight their respective impacts on model performance.
@@
(d) ImageNet-R-$\eta$ Tuning
(e) ImageNet-R-$\upsilon$ Tuning
(f) ImageNet-R-$\lambda$ Tuning
Figure 5: Illustration of the tuning process for each hyper-parameter on CIFAR100 and ImageNet-R, showing the continuous performance variations.

### v2

@@ 5.1 Extend Sample Size Through PEFT
Drawing upon the theoretical underpinnings elucidated in lemma 3 and [2], it becomes evident that the augmentation of task-specific sample size exerts a significant influence on mitigating generalization discrepancies. In light of this insight, we introduce a novel strategy meticulously tailored for the PEFT paradigm, predicated on the existence of an optimal function $f_{0}^{*}(x)$, as rigorously defined in Eq. 3. This approach operates across three specialized subnetworks, each responsible for feature generation within unique representational space, thereby engendering a composite feature set. This process not only amplifies the effective sample (feature) size pertinent to each subtask but also fosters a more nuanced and comprehensive representation of the underlying data manifold. Through the judicious adjustment of subnetwork parameters $p_{i}$, facilitated by the integration of these multi-dimensional feature representations, our proposed framework achieves a tripling of the representational scope for individual samples. More importantly, we can replace different types of subnetworks to enable the model to adaptively learn the same image in different representational spaces, thereby avoiding the need for human-provided prior processing [105, 41, 16, 17] at the image level and reducing additional optimization overhead. This enhancement is systematically illustrated through the intricate adaptive interactions depicted in Fig. 2.
@@
PEFT-CL methodologies typically employ a prompt pool or introduce auxiliary parameters while preserving pre-trained weights, modifying $E_{1}^{i},E_{2}^{i},\ldots,E_{N}^{i}$ within each transformer block to influence the class token $E_{CLS}$. This generates a novel feature space that adapts to subtasks and mitigates catastrophic forgetting. In these methods, the predetermined task prompt pool is traditionally used to derive task-specific embeddings, selecting prompts through cosine similarity [88, 87, 39]. While effective, this paradigm incurs substantial computational overhead when intervening in the self-attention mechanism and constrains the network’s capacity for generating diverse, instance-specific adaptive interventions dynamically. To address these limitations, our proposed NTK-CL framework implements a more efficient paradigm utilizing additional trainable parameters to autonomously generates instance-specific interventions. These interventions then interact with our proposed feature space post-multi-head self-attention (MSA) module to yield task-specific embeddings. This approach not only maximizes the utilization of pre-trained knowledge but also effectively reduces the computational burden brought by intervening MSA calculations.
@@
Figure 3: The illustration depicts the fusion of multi-level features to generate three distinct features per sample, thereby increasing the sample size available for model optimization.
@@ 5.2 Task-Level Feature Constraints
Therefore, we propose a compromise approach. Within the context of NTK theory, the optimization of infinitely wide neural networks mirrors a Gaussian process [11, 44], yielding a locally constant NTK matrix [36, 46, 15]. Given this, it is reasonable to assume that $\Phi^{*}(X_{\tau},X_{k})=\Phi_{0}(X_{\tau},X_{k})=\Phi_{1}(X_{\tau},X_{k})=\cdots=\Phi_{\infty}(X_{\tau},X_{k})$. Moreover, networks pre-trained on extensive datasets emulate the properties of infinitely wide networks [45, 89, 80], aligning with our pre-trained model. Therefore, we relax the original constraint, assuming that the pre-trained model is at this local optimum.
@@
Knowledge Retention: Achieving the retention of past knowledge is a critical component in traditional CL methods [8, 48, 49]. However, in PEFT-CL methods, this fundamental aspect has been notably underemphasized. Contemporary PEFT-CL methods predominantly involve repositioning prompts [88, 25] or jointly utilizing all task-specific subnetwork parameters [51, 109], thereby shifting the focus away from the retention of past knowledge to the training performance of each task. However, such strategies necessitate the maintenance of optimal parameter configurations for each encountered task, which not only incurs substantial storage demands but also potentially limits the system’s adaptability, particularly in scenarios characterized by a high density of tasks. To mitigate these challenges, we propose a paradigm shift that emphasizes the reevaluation of knowledge retention mechanism, eschewing the necessity for per-task parameter storage. Central to our method is the introduction of an adaptive Exponential Moving Average (EMA) mechanism. This mechanism, as depicted in Fig. 4, facilitates a more streamlined and scalable solution to the catastrophic forgetting problem, enhancing the overall efficiency and efficacy of PEFT-CL systems.
@@
Traditional EMA applications often maintain a static base model, incrementally integrating optimized weights to preserve historical data. However, this approach proves suboptimal in PEFT-CL settings due to the substantial disparities in weights across tasks. Directly preserving a large proportion of past weights can detrimentally affect the performance on current task, while retaining an entire model’s weights is excessively redundant. Therefore, we propose two improvements. First, we categorize the adaptation parameters responsible for generating embedding into two segments: $p^{pre}$ for historical knowledge and $p^{curr}$ for current insights. Secondly, we apply the EMA mechanism exclusively to the adaptation modules’ parameters, leaving other optimizable parameters untouched to ensure the optimization remains streamlined. Throughout the optimization of task $\tau+1$, only $p^{curr}_{\tau+1}$ is modified, while $p^{pre}_{\tau+1}$ is adaptively adjusted post-task-$\tau$ completion, employing an adaptive EMA scheme:
@@
Under this mechanism, each past task equitably contributes to constructing embeddings related to historical knowledge without compromising the current task’s insights, while avoiding the excessive memory overhead of storing parameters for each task, as seen in [109]. Consequently, $E_{CLS}^{S1}$, $E_{CLS}^{S2}$, and $E_{CLS}^{HAE}$ all consist of two components: $concat[f(x,p_{pre}),f(x,p_{curr})]$.
Task-Feature Dissimilarity:  [^3] Regarding why task-feature orthogonality does not impair the propagation and retention of knowledge among similar classes across different tasks, we provide further explanations in Appendix G. Based on the findings in [76, 66], it is evident that achieving class-level orthogonal insulation can effectively enhance the performance of PEFT-CL models. However, our theoretical analysis in Section 5.2 and insights from [20, 4] indicate that achieving task-level orthogonal insulation between $f^{*}(X_{\tau})$ and $f^{*}(X_{k})$ is sufficient to reduce the generalization gap and obtain good continual performance. This task-level orthogonal insulation not only simplifies the model requirements but also ensures robust and efficient learning across tasks. Therefore, relying on the prototype classifier, we propose an optimization loss. In line with [109, 82], we update the prototype classifier $\zeta$ upon completion of each task’s optimization and strictly prohibit accessing previous samples in subsequent optimizations to comply with PEFT-CL constraints. During the optimization of task $\tau$, we randomly sample $\zeta_{\tau}$ from $\zeta$ to represent $f^{*}(X_{\tau})$  [^4] Sampling from the parameter space of the prototype classifier $\zeta$, unlike approaches such as Hide-Prompt [82] and APG [78], avoids compressing past embedding distributions and adding extra training overhead. This method also eliminates the need for a replay buffer, effectively bypassing the typical constraints associated with PEFT-CL.. To initially distinguish $f^{*}(X_{\tau})$ from $f^{*}(X_{k})$, we use the InfoNCE [60] as a metric, employing $\zeta_{\tau}$ as the negative sample, while using samples $x_{\tau}$ (represented by $E_{CLS}^{HAE}$, as this is the feature used for final classification) from task $\tau$ as positive samples.
@@
To further ensure orthogonality between $f^{*}(X_{\tau})$ and $f^{*}(X_{k})$, we apply the truncated SVD method [26] to constrain the optimization of $f^{*}(X_{k})$. Specifically, we decompose $\zeta$ to obtain the orthogonal basis $\mathbf{U}$ that defines the classification (preceding feature) space. We then map $x_{\tau}$ into this space and remove the unmappable part from the original $x_{\tau}$. When the retained mappable portion is sufficiently small, the orthogonality between $x_{\tau}$ and $\zeta$ is ensured.
@@
where $\tilde{proj}(a,b)$ represents the unmappable portion of $a$ within the space spanned by the orthogonal basis functions decomposed from $b$.
@@ 5.3 Regularization Adjustment
In accordance with the theoretical constraints delineated in Appendix B, which advocate for the incorporation of ridge regression to ensure a well-conditioned solution, we deploy an L2 regularization. As specified in Eq. 32, the regularization term is structured as $\left\|p_{\tau}-p_{\tau-1}^{*}\right\|_{2}^{2}$, targeting the parameter shifts from task $\tau-1$ to task $\tau$. Consequently, we meticulously design our regularization term to mirror this structure and temporarily retain the trainable parameters $p^{pre}$ from the preceding task. This targeted regularization is then precisely applied to the parameters of the various modules within our NTK-CL, formulated as follows:
@@
$\mathcal{L}_{total}=\mathcal{L}_{cls}+\eta\mathcal{L}_{dis}+\upsilon\mathcal{L}_{orth}+\lambda\mathcal{L}_{reg},$ (29)
@@
Testing Evaluation: Upon concluding the training regimen, the evaluation phase commences with simultaneous testing across all tasks. This phase distinctly prioritizes the synthesized hybrid adaptation features $E_{CLS}^{HAE}$ for final analysis. Through the final prototype classifier $\zeta$, these features are transformed into logits, which are aligned with the corresponding labels to deduce the test accuracy.
Algorithm 1 NTK-CL Framework for PEFT-CL
1:Pre-trained model $f^{*}_{0}$; task set $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$; initial PEFT parameters $p_{1}=p_{1}^{\text{pre}}\oplus p_{1}^{\text{curr}}$, where $p_{1}^{\text{pre}}=p_{1,S1}^{\text{pre}}\oplus p_{1,S2}^{\text{pre}}\oplus p_{1,HAE}^{\text{pre}}$ and $p_{1}^{\text{curr}}=p_{1,S1}^{\text{curr}}\oplus p_{1,S2}^{\text{curr}}\oplus p_{1,HAE}^{\text{curr}}$; hyper-parameters $\eta$, $\upsilon$, $\lambda$; learning rate $\xi$
2:Trained PEFT-CL model $f^{*}_{T}=f^{*}_{0}\circ p^{*}_{T}$ with minimized generalization error
3: Initialize frozen $p_{1}^{\text{pre}}$ and trainable $p_{1}^{\text{curr}}$
4:for each task $\mathcal{D}_{\tau}$ in $\mathcal{D}$ do
5:    Retrieve task-specific data $(X_{\tau},Y_{\tau})$
6:    Compute features:
7:$E_{CLS}^{S1}=(f^{*}_{0}\circ p_{\tau,S1}^{\text{pre}})(X_{\tau})\oplus(f^{*}_{0}\circ p_{\tau,S1}^{\text{curr}})(X_{\tau})$
8:$E_{CLS}^{S2}=(f^{*}_{0}\circ p_{\tau,S2}^{\text{pre}})(X_{\tau})\oplus(f^{*}_{0}\circ p_{\tau,S2}^{\text{curr}})(X_{\tau})$
9:$E_{CLS}^{HAE}=p_{\tau,HAE}^{\text{pre}}(E_{CLS}^{S1},E_{CLS}^{S2})\oplus p_{\tau,HAE}^{\text{curr}}(E_{CLS}^{S1},E_{CLS}^{S2})$
10:    Compute the classification loss $\mathcal{L}_{\text{cls}}$ as in Eq. 23:
11:$\mathcal{L}_{\text{cls}}=CE(E_{CLS}^{S1},Y_{\tau})+CE(E_{CLS}^{S2},Y_{\tau})+CE(E_{CLS}^{HAE},Y_{\tau})$
12:    Enforce task-level orthogonality constraints $\mathcal{L}_{\text{dis}}$ and $\mathcal{L}_{\text{orth}}$ for $E_{CLS}^{HAE}$ according to Eq. 26 and Eq. 27
13:    Apply parameter regularization $\mathcal{L}_{\text{reg}}$ as per Eq. 28:
14:$\mathcal{L}_{\text{reg}}=\sum_{i\in\{S1,S2,HAE\}}\|p_{\tau,i}^{\text{curr}}-p_{\tau,i}^{\text{pre}}\|_{2}^{2}$
15:    Compute the overall loss using Eq. 29:
16:$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cls}}+\eta\mathcal{L}_{\text{dis}}+\upsilon\mathcal{L}_{\text{orth}}+\lambda\mathcal{L}_{\text{reg}}$
17:    Update $p_{\tau}^{\text{curr}}$ using backpropagation optimization:
18:$p_{\tau}^{\text{curr}}\leftarrow p_{\tau}^{\text{curr}}-\xi\nabla_{p_{\tau}^{\text{curr}}}\mathcal{L}_{\text{total}}$
19:    Update the prototype classifier with $E_{\text{CLS}}^{\text{HAE}}$
20:   if $\tau$ is not the last task then
21:      Perform Adaptive EMA updates on $p^{\text{pre}}_{\tau}$ as per Eq. 24 and Eq. 25
22:   end if
23:end for
24: return Final model $f^{*}_{T}$
In summary, the operational sequence of our NTK-CL framework is encapsulated in Algorithm 1.
@@ 6 Experiments
In this study, we utilize a carefully curated suite of benchmark datasets designed to support a rigorous and comprehensive evaluation of model generalization within the PEFT-CL paradigm. These datasets encompass a wide spectrum of domains, including general object recognition, domain adaptation, fine-grained animal classification, earth observation, agricultural analytics, task adaptation, and healthcare diagnostics. This diverse selection ensures a robust evaluation framework that captures the complexities and challenges inherent in real-world applications. Detailed descriptions of each dataset, along with the corresponding training protocols, evaluation metrics, and implementation specifics, are provided in Table I and Appendix E, thereby promoting reproducibility and facilitating a clearer understanding.
@@
TABLE VI: Ablation study on the ViT-B/16-IN21K model, evaluating its performance across the CIFAR100 and ImageNet-R datasets. The study features a detailed breakdown of model components, denoted in each column by the inclusion (✓) of specific modules and strategies: Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid), Knowledge Retention (KR), Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). Incremental accuracies ($\bar{A}$) are reported to highlight their respective impacts on performance.
@@
(d) CIFAR100 AHPS
(e) ImageNet-R-$\eta$ Tuning
(f) ImageNet-R-$\upsilon$ Tuning
(g) ImageNet-R-$\lambda$ Tuning
(h) ImageNet-R AHPS
Figure 5: Performance comparison of NTK-CL with different hyper-parameter settings and the proposed Automatic Hyper-Parameter Search (AHPS) strategies, based on Bayesian optimization or dynamic loss scaling, on CIFAR-100 and ImageNet-R.